OpenStudy (anonymous):

5x – 7(x + 1) > –9 + 3

Elsa213 (elsa213):

" Simplifying 5x + -7(x + 1) = -9 + 3 Reorder the terms: 5x + -7(1 + x) = -9 + 3 5x + (1 * -7 + x * -7) = -9 + 3 5x + (-7 + -7x) = -9 + 3 Reorder the terms: -7 + 5x + -7x = -9 + 3 Combine like terms: 5x + -7x = -2x -7 + -2x = -9 + 3 Combine like terms: -9 + 3 = -6 -7 + -2x = -6 Solving -7 + -2x = -6 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '7' to each side of the equation. -7 + 7 + -2x = -6 + 7 Combine like terms: -7 + 7 = 0 0 + -2x = -6 + 7 -2x = -6 + 7 Combine like terms: -6 + 7 = 1 -2x = 1 Divide each side by '-2'. x = -0.5 Simplifying x = -0.5 "

Elsa213 (elsa213):

\(\rlap{\rlap{\rlap{\sf\LARGE\color{greenyellow}{~Welcome~~~To~~~~OpenStudy}}}}\) \(\rlap{\rlap{\rlap{\sf\LARGE\color{#00cfff}{Welcome~To~~OpenStudy}}}}\) \(\rlap{\rlap{\rlap{\cal\tiny\color{teal}{Please~Read~The~Code~Of~Conduct~When~You~Get~The~Chance.}}}}\) \(\rlap{\rlap{\rlap{\sf\Large\color{black}{~~~~~~Thank~You\heartsuit}}}}\) \(\rlap{\rlap{\rlap{\sf\LARGE\color{aquamarine}{~Thank~}\color{#00cfff}{~You}}}}\)